OpenStudy (anonymous):

Explain why #4 is A.

OpenStudy (anonymous):

I have no idea...where is #4

OpenStudy (anonymous):

\[\sin ^{-1} \sqrt{3}/2\]

OpenStudy (anonymous):

the inverse sin of a value is a measurement of an angle so the sin of some angle = \[\frac{\sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

\[\sin \frac{ \pi }{ 3 } = \frac{ \sqrt{3} }{ 2 } \] this will occur a full revolution from \[\frac{ \pi }{ 3 }\] which will happen at \[2 \pi from \frac{ \pi }{ 3 }\] \[2 \pi * n \] will be n revolutions around the circle and always land on \[\frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

I hope that makes sense...